# A new approach to principal–agent problems with volatility control

### v1

@@ 3.3 Penalisation contracts
    Recall that, in 2, for a given couple $(Z,\Gamma)\in{\cal V}$, the agent’s optimal efforts are given by the maximiser of the Hamiltonian ${\cal H}_{\rm A}$, namely $\nu^{\star}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$ $t\in[0,T]$, where the function $u^{\star}$ is defined by (2.13). On the other hand, in 4, for $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$, the agent’s optimal efforts are given by the maximiser of the (constrained) Hamiltonian ${\cal H}^{\circ}_{\rm A}$, namely $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})$, $t\in[0,T]$, where the function $u^{\circ}$ is defined by (3.19). Therefore, the first step is to prove that for any $(Z,\Gamma)\in{\cal V}$, there exist $\Sigma\in{\cal S}$ such that $Z\in{\cal V}^{\circ}(\Sigma)$ and satisfying
    $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})=:\nu^{\star}_{t},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$ (3.21)
    To have a one-to-one correspondence, one need to prove that the converse also holds, namely that for any $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$, there exist $\Gamma$ such that $(Z,\Gamma)\in{\cal V}$ and (3.21) holds $\mathrm{d}t\otimes\mathbb{P}^{\circ}$–a.e. on $[0,T]\times\Omega$.
Step 1. To prove the first step, let $(Z,\Gamma)\in{\cal V}$. Note that, by definition of the function $u^{\star}$, we have:
${\cal H}_{\rm A}(t,x,y,z,\gamma)=h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)},\quad(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}.$
Alternatively, by Remark 3.5, we have for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$,
${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $=\sup_{S\in\mathbb{S}_{t}(x)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)+\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$
Recall the definition of the function $u^{\circ}$ in (3.19) as the maximiser of the (constrained) Hamiltonian ${\cal H}^{\circ}_{\rm A}$, namely
$u^{\circ}(t,x,y,z,S)\in\operatorname*{arg\,max}_{u\in U_{t}^{\circ}(x,S)}h^{\circ}_{\rm A}(t,x,y,z,u),\quad(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}.$
Therefore, for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have by definition of the set $U_{t}^{\circ}(x,S)$,
${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $=\sup_{S\in\mathbb{S}_{t}(x)}\bigg{\{}h^{\circ}_{\rm A}(t,x,y,z,u^{\circ}(t,x,z,y,S))+\dfrac{1}{2}{\rm Tr}\big{[}\gamma[\sigma(\sigma)^{\top}](t,x,u^{\circ}(t,x,z,y,S))\big{]}\bigg{\}}$
$=\sup_{S\in\mathbb{S}_{t}(x)}\big{\{}h_{\rm A}(t,x,y,z,\gamma,u^{\circ}(t,x,z,y,S))\big{\}},$
where we also used the definition of $h_{\rm A}$ in (2.10). Hence, for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have
$h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)}$ $=\sup_{S\in\mathbb{S}_{t}(x)}\big{\{}h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\circ}(t,x,z,y,S)\big{)}\big{\}},$
and we conclude that there exists a map $S^{\star}$ such that for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have $S^{\star}(t,x,y,z,\gamma)\in\mathbb{S}_{t}(x)$ and
@@
$\Sigma_{t}:=S^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})\in\operatorname*{arg\,max}_{S\in\mathbb{S}_{t}(X)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},S)+\dfrac{1}{2}{\rm Tr}\big{[}\Gamma_{t}S\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$
@@
Step 2. To prove that the converse also holds, let $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$. We basically want to show that there exist $\Gamma$ such that $(Z,\Gamma)\in{\cal V}$ and (3.21) holds. For this, we introduce the dual problem of ${\cal H}^{\circ}_{\rm A}$, namely
~$\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S):=\inf_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}},\quad(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}.$ (3.22)
First, using again the link (3.11) between both Hamiltonians ${\cal H}_{\rm A}$ and ${\cal H}^{\circ}_{\rm A}$, it is clear that
@@
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)\leq{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]},\quad(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d},\;S\in\mathbb{S}_{t}(x),$
and thus, taking the infimum over $\gamma\in\mathbb{M}^{d}$,
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)\leq\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S).$ (3.23)
Furthermore, as already noticed in [9, Remark 3.5], the Hamiltonian ${\cal H}_{\rm A}$ is convex in $\gamma\in\mathbb{M}^{d}$, implying here that the dual problem introduced above by (3.22) is a convex minimisation problem. Moreover, since there exists an optimal solution for the primal problem and the dual variable is unconstrained, there exists an optimal solution for the dual problem, i.e. a minimiser $\gamma^{\star}:=g^{\star}(t,x,y,z,S)\in\mathbb{M}^{d}$. By definition of $u^{\star}$ and $h_{\rm A}$, we thus have
~$\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=\inf_{\gamma\in\mathbb{M}^{d}}\bigg{\{}h_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma)\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}=h_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma^{\star})\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma^{\star}S\big{]},$
for all $(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}$. Moreover, since the dual problem is convex in $\gamma$, we have
$[\sigma\sigma^{\top}](t,x,u^{\star}(t,x,y,z,\gamma))-S\in 2\partial_{\gamma}\bigg{(}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{)},$
where here, $\partial_{\gamma}$ denotes the subgradient with respect to $\gamma$. Since $\gamma^{\star}$ is a minimiser, we necessarily have
$[\sigma(\sigma)^{\top}]\big{(}t,x,u^{\star}(t,x,y,z,\gamma^{\star})\big{)}=S,$ (3.24)
for all $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$. We thus deduce, using the definitions of $h^{\circ}_{\rm A}$ and ${\cal H}_{\rm A}^{\circ}$ that
~$\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma^{\star})\big{)}\leq\sup_{u\in U_{t}^{\circ}(x,S)}h^{\circ}_{\rm A}\big{(}t,x,y,z,u\big{)}={\cal H}^{\circ}_{\rm A}(t,x,y,z,S),$
since $u^{\star}(t,x,y,z,\gamma^{\star})\in U^{\circ}_{t}(x,S)$ by (3.24). Combining with (3.23), we conclude that
$h^{\circ}_{\rm A}\big{(}t,x,y,z,S,u^{\circ}(t,x,y,z,S)\big{)}={\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma^{\star})\big{)},$
for all $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$ and $S\in\mathbb{S}_{t}(x)$, implying in particular
$u^{\circ}(t,x,y,z,S)=u^{\star}(t,x,y,z,\gamma^{\star}),\;\text{ for }\;\gamma^{\star}\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$
In other words, for any $\Sigma\in{\cal S},Z\in{\cal V}^{\circ}(\Sigma)$ and the corresponding state variables $(X,Y)$ with dynamics given by (3.20), one can define $\Gamma$ as follows,
@@
$\Gamma_{t}:=g^{\star}(t,X,Y_{t},Z_{t},\Sigma_{t})\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,X,Y_{t},Z_{t},\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma\Sigma_{t}\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{-a.e. on }[0,T]\times\Omega.$
Moreover, the integrability conditions in Definition 2.4 follows from the integrability conditions in Definition 3.6, implying that $(Z,\Gamma)\in{\cal V}$, and by the previous reasoning, the desired equality of the optimal efforts. ∎
@@
To summarise, using Lemma 2.5, Lemma 3.11, Theorem 3.9 and finally Lemma 3.4, we obtain the following relationship for the values of the various problems,
@@
which directly implies equality of all the values, thus concluding the proof of Theorem 3.10. In other words, the four problems outlined are in fact equivalent. While 1 and 3 cannot generally be solved directly, 2 and 4 correspond to more standard stochastic control problems, and should therefore be easier to solve, through standard techniques relying on Hamilton-Jacobi-Bellman equations. We illustrate in the following section the link between 2 and 4 and their solutions through a couple of toy examples with explicit solutions. In particular, these examples will allow to better highlight the link between the controls $\Gamma$ in 2 and $\Sigma$ in 4, which is not quite explicit in the above result Lemma 3.11, because of the general formulation.

### v2

@@ 3.3 Penalisation contracts
    Fix $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$. First, by definition of the function $u^{\star}$ in (2.13),
    ${\cal H}_{\rm A}(t,x,y,z,\gamma)=h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)}.$
    Alternatively, by Remark 3.5, we have
    ${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $=\sup_{S\in\mathbb{S}_{t}(x)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)+\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}},$
    recalling that $\mathbb{S}_{t}(x)$ is defined for all $(t,x)\in[0,T]\times\Omega$ by (3.9). Define $S^{\star}:=\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,x,u^{\star}(t,x,y,z,\gamma)\big{)}$. By definition, $S^{\star}\in\mathbb{S}_{t}(x)$, and we thus have
    ${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $\geq{\cal H}^{\circ}_{\rm A}\big{(}t,x,y,z,S^{\star}\big{)}+\dfrac{1}{2}{\rm Tr}\big{[}\gamma\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,x,u^{\star}(t,x,y,z,\gamma)\big{)}\big{]}.$
    Moreover, also by definition of $S^{\star}$, we have $u^{\star}(t,x,y,z,\gamma)\in U_{t}^{\circ}(x,S^{\star})$, implying
    ${\cal H}^{\circ}_{\rm A}\big{(}t,x,y,z,S^{\star}\big{)}:=\sup_{u\in U_{t}^{\circ}(x,S^{\star})}\;h^{\circ}_{\rm A}(t,x,y,z,u)\geq h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma)\big{)}.$
    Combining all the previous equations we obtain the following inequalities,
    $h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)}$ $\geq{\cal H}^{\circ}_{\rm A}\big{(}t,x,y,z,S^{\star}\big{)}+\dfrac{1}{2}{\rm Tr}\big{[}\gamma\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,x,u^{\star}(t,x,y,z,\gamma)\big{)}\big{]}$
    $\geq h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma)\big{)}+\dfrac{1}{2}{\rm Tr}\big{[}\gamma\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,x,u^{\star}(t,x,y,z,\gamma)\big{)}\big{]}$
    $=h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)},$
    which necessarily implies equality at every lines, and in particular
    ${\cal H}^{\circ}_{\rm A}\big{(}t,x,y,z,S^{\star}\big{)}=h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma)\big{)}.$
    Therefore, the map $u^{\star}$ is a maximiser of the constrained Hamiltonian ${\cal H}^{\circ}_{\rm A}$ for the quadratic variation $S^{\star}$.
To summarise, there exists a map $S^{\star}$ such that for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have $S^{\star}(t,x,y,z,\gamma)\in\mathbb{S}_{t}(x)$ and
@@
$\Sigma_{t}:=S^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})=\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,X,u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})\big{)},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega,$
@@
The previous lemma highlights that, for any $y_{0}\geq R_{\rm A}$ and $(Z,\Gamma)\in{\cal V}$, one can find an appropriate process $\Sigma\in{\cal S}$ such that $Z\in{\cal V}^{\circ}(\Sigma)$, and the optimal efforts $\nu^{\star}$ and $\nu^{\circ}$ coincide. Given the formulation of Problem 2 and Problem 4, this is enough to conclude that $\widetilde{V}_{\rm P}\geq\widetilde{V}^{\circ}_{\rm P}$. We now turn to the converse inequality, for which we require the following assumption.
Assumption B.
For all $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$ and $S\in\mathbb{S}_{t}(x)$, we have
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=\inf_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$ (3.22)
Moreover, the infimum is achieved, namely there exists a measurable function $\gamma^{\star}:[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{d}^{+}\longrightarrow\mathbb{M}^{d}$ such that
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)={\cal H}_{\rm A}\big{(}t,x,y,z,\gamma^{\star}(t,x,y,z,S)\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma^{\star}(t,x,y,z,S)S\big{]}.$ (3.23)
Remark 3.12.
While the previous assumption is not explicitly stated in [9], the proof of the main result regarding the optimality of the contract form (2.11) may fail if this assumption is not satisfied. More precisely, it is claimed in this proof that the function defined in (4.17), namely using our notations
@@
$\widebar{f}:\gamma\in\mathbb{M}^{d}\longmapsto{\cal H}_{\rm A}(t,x,y,z,\gamma)-{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]},\quad(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{d}^{+},$
is surjective on $(0,\infty)$. However, for this claim to be true, one need at least
$\inf_{\gamma\in\mathbb{M}^{d}}\widebar{f}(\gamma)=0,$
and this previous condition can be rewritten as (3.22), for all $(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{d}^{+}$. In other words, $(-2\,{\cal H}^{\circ}_{\rm A})$ should be the convex conjugate of $2\,{\cal H}_{\rm A}$. However, as already mentioned in Remark 3.5, $2\,{\cal H}_{\rm A}$ is already the convex conjugate of $(-2\,{\cal H}^{\circ}_{\rm A})$. Therefore, the previous condition is true if and only if $(-2\,{\cal H}^{\circ}_{\rm A})$ actually coincide with its biconjugate (the convex conjugate of the convex conjugate). Note that, for any function $g$, its biconjugate $g^{\star\star}$ satisfies $g^{\star\star}\leq g$, which implies here $(-2\,{\cal H}^{\circ}_{\rm A})^{\star\star}\leq-2\,{\cal H}^{\circ}_{\rm A}$, and thus
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)\leq\inf_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$
However, by the Fenchel–Moreau theorem, the inequality $g^{\star\star}\leq g$ becomes an equality if and only if $g$ is convex, lower semi-continuous and proper. Since ${\cal H}^{\circ}_{\rm A}$, as a function of $S$, does not always satisfy these conditions in the framework under consideration here and in [9], this naturally leads to the first part of Assumption B. The second part of the assumption, namely that the infimum is achieved, is also needed in the proof of [9, Theorem 3.6]. More precisely, from the fact that $\widebar{f}$ is surjective on $(0,\infty)$, it is deduced that for any $\dot{K}\geq 0$, there exists a measurable process $\Gamma$ such that
$\dot{K}_{t}={\cal H}_{\rm A}(t,X,Y_{t},Z_{t},\Gamma_{t})-{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-\dfrac{1}{2}{\rm Tr}\big{[}\Gamma_{t}\Sigma_{t}\big{]}.$
However, as mentioned, it is possible that $\dot{K}_{t}=0$ for some $t\in[0,T]$, and thus one need the function $\widebar{f}$ to actually be surjective onto $[0,\infty)$ instead of $(0,+\infty)$ only. This naturally leads to (3.23) in the second part of Assumption B. [^2] We have been in contact with the authors of [9] to verify the relevance of our additional assumption. We are currently discussing potential ways to construct alternative forms of contracts in settings where Assumption B does not hold.
Under Assumption B, we can state the following result, which serves as the counterpart to Lemma 3.11. Its proof follows similar arguments, adapted to establish the reverse inequality, and is reported below for completeness.
Lemma 3.13.
Let $y_{0}\geq R_{\rm A}$, $\Sigma\in{\cal S}$, $Z\in{\cal V}^{\circ}(\Sigma)$, and consider the corresponding state variables $(X,Y)$ with dynamics given by (3.21), as well as an optimal effort $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})$, $t\in[0,T]$. Under Assumption B, there exists a process $\Gamma\in\mathbb{M}^{d}$ defined by
@@
$\Gamma_{t}\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,X,Y_{t},Z_{t},\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma\Sigma_{t}\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{-a.e. on }[0,T]\times\Omega,$
such that $(Z,\Gamma)\in{\cal V}$ and $\nu^{\star}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})$, $\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{-a.e. on }[0,T]\times\Omega$.
  • Proof of Lemma 3.13.
    Fix $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$ and $S\in\mathbb{S}_{t}(x)$. On the one hand, by definition of the function $u^{\circ}$ in (3.20), we have
    ${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=h_{\rm A}^{\circ}\big{(}t,x,y,z,u^{\circ}(t,x,y,z,S)\big{)}.$
    On the other hand, by Assumption B, there exists a measurable function $\gamma^{\star}:[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{d}^{+}\longrightarrow\mathbb{M}^{d}$ such that (3.23) holds, namely
    ${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)={\cal H}_{\rm A}\big{(}t,x,y,z,\gamma^{\star}(t,x,y,z,S)\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma^{\star}(t,x,y,z,S)S\big{]}.$
    Then, by definition of ${\cal H}_{\rm A}$, we have for any $\gamma\in\mathbb{M}^{d}$
    ${\cal H}_{\rm A}(t,x,y,z,\gamma)\geq h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\circ}(t,x,y,z,S)\big{)},$ (3.24)
    and since $[\sigma\sigma^{\top}](t,x,u^{\circ}(t,x,y,z,S))=S$, we deduce
    ${\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}$ $\geq h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\circ}(t,x,y,z,S)\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma[\sigma\sigma^{\top}](t,x,u^{\circ}(t,x,y,z,S))\big{]}$
    $=h_{\rm A}^{\circ}\big{(}t,x,y,z,u^{\circ}(t,x,y,z,S)\big{)}.$
    Combining all the previous equalities and inequalities together for $\gamma:=\gamma^{\star}(t,x,y,z,S)$, we have
    $h_{\rm A}^{\circ}\big{(}t,x,y,z,u^{\circ}(t,x,y,z,S)\big{)}$ $={\cal H}^{\circ}_{\rm A}(t,x,y,z,S)\geq h_{\rm A}^{\circ}\big{(}t,x,y,z,u^{\circ}(t,x,y,z,S)\big{)},$
    thus implying equality everywhere, in particular in (3.24) for $\gamma:=\gamma^{\star}(t,x,y,z,S)$. Therefore, the map $u^{\circ}$ is a maximiser of the Hamiltonian ${\cal H}_{\rm A}$ when choosing $\gamma^{\star}$.
To summarise, there exists a map $\gamma^{\star}$ such that for all $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$ and $S\in\mathbb{S}_{t}(x)$, we have
$u^{\star}\big{(}t,x,y,z,\gamma^{\star}(t,x,y,z,S)\big{)}=u^{\circ}(t,x,y,z,S).$
In other words, for any $\Sigma\in{\cal S}$, $Z\in{\cal V}^{\circ}(\Sigma)$ and the corresponding state variables $(X,Y)$ with dynamics given by (3.21), one can define a process $\Gamma$ by taking
$\Gamma_{t}=\gamma^{\star}(t,X,Y_{t},Z_{t},\Sigma_{t}),\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega$
It is then straightforward to verify that $(Z,\Gamma)\in{\cal V}$ and, by the previous reasoning,
$\nu^{\star}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t}),\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$
Using Lemma 3.11 and Lemma 3.13, we naturally deduce the following proposition.
Proposition 3.14.
Under Assumption B, solving Problem 2 is equivalent to solving Problem 4, i.e. $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$.
@@
To summarise, using sequentially Lemma 2.5, Proposition 3.14, Theorem 3.9 and finally Lemma 3.4, we obtain the following relationship for the values of the various problems,
@@
which directly implies equality of all the values, and in particular concludes the proof of Theorem 3.10. In other words, under Assumptions A and B, the four problems outlined are in fact equivalent. We recall here that Assumption A is a standing assumption, required to be able to derive a ‘simple’ form of contracts in the ‘first-best’ problem, as mentioned in Remark 3.3. This assumption is not required to develop the approach in [9]. On the other hand, although Assumption B is not stated explicitly in [9], it is required to establish the equivalence between the two problems and ensure optimality of the contracts derived in [9], as explained in Remark 3.12. While this assumption is satisfied in most of the models studied in the literature, see for example Sections 4.1 and 4.2, it may fail in some frameworks. In particular, Section 4.3 presents a ‘counter-example’, in which Assumption B is not satisfied.